OpenStudy (anonymous):

Solve the equation step by step. List each step, along with the property you used during that step. –5 × 1 × 11 × 4.

OpenStudy (nurali):

−5×1×11×4 =−5×44 =−220

OpenStudy (nurali):

the properties of multiplication

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

Summarize the order of operations for evaluating expressions. Use complete sentences.

OpenStudy (nurali):

It can be summarized using the PEMDAS rule. It stands for: P ==> Parentheses first E ==> Exponents second MD ==> Multiplication and Division AS ==> Addition and Subtraction

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

Anytime.